### v1

@@ 4.1 Partition Optimization and its Hardness
As illustrated above, the choice of the item partition plays an important role in both the performance and robustness of Mechanism 2. Below, we formalize two different metrics for partitions, which determine the performance and robustness of the resulting mechanism, respectively. We will study their implications for the efficient computation of a high-quality partition, as well as their intrinsic tradeoff.
  • Informativeness (as optimization objective). In regard to the performance, we aim to identify partitions that are most informative for the review score calibration. As mentioned previously, generally larger-sized blocks are preferred — in the extreme case, if a block contains only one item, the mechanism cannot elicit any ranking information on this single item in the block. However, to get large blocks, it may sometimes be unavoidable to have a small amount of small or even trivial blocks. This creates a complex tradeoff about partition choices. To systemically evaluate the coverage quality of a partition ${\mathcal{S}}$, we introduce an informativeness objective function
    $\operatorname{obj}({\mathcal{S}})=\sum_{k=1}^{K}w(|{\mathcal{S}}_{k}|),$
    determined by any weight function $w$ on the size of each partition block. For example, two choices of weight function $w(\cdot)$ are of particular interest. One is $w(x)=x^{2}$ as a simple quadratic function of block size, which relates to the number of pairwise comparisons within a block:
@@
    Another natural choice is $w(x)=\mathop{max}\{x-1,0\}$, which is called size-based defined as follows:
@@
    where the second equation is because any block in the partition has at least $1$ item (i.e., $|{\mathcal{S}}_{k}|\geq 1$). Since $n$ is a constant, maximizing the size-based objective is equivalent to minimizing the size of the partition $K$ (thus the name “size-based”).
    Generally, it is reasonable to expect that function $w(\cdot)$ should satisfy increasing monotonicity and convexity — i.e., the larger the size is, the faster the objective increases, and the better the Isotonic Mechanism performs. Another natural requirement is $w(0)=0$ in order to prevent empty blocks in the optimal partitions. Both functions above satisfy these properties, and so do any $w(x)=x^{p}$ for $p\geq 1$. However, it appears to be an intractable task to precisely describe the relationship between the structure of partition and the statistical guarantee of the mechanism. Thus, in Section 4.2, we seek to overcome this challenge by designing algorithms that can simultaneously perform well for many natural weight functions satisfying these properties.
@@
  • Robustness (as optimization constraint). If the informativeness metric above is viewed as the objective we want the partition to maximize, then the robustness metric can be viewed as a constraint we wish the partition to satisfy. In regard to robustness, aggregating information from multiple owners helps to mitigate each individual owner’s noisy perception of their items. Thus, the minimal number of owners for each partition block matters. Formally, we say a partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ is $L$-strong if there are at least $L$ owners who share all items in the same partition block, i.e., $\left|{\mathcal{T}}_{k}\right|\geq L,\forall{\mathcal{S}}_{k}\in{\mathcal{S}}$ with $|{\mathcal{S}}_{k}|>1$. The larger $L$ is, the less prone the Isotonic Mechanism is to each individual owner’s randomness in her perception. Notably, $L\geq 1$ is the bare minimum requirement for Mechanism 2 to work. It is intuitive that robustness and informativeness are a tradeoff in general: a larger $L$ necessarily leads to less informative partitions, since a stricter constraint necessarily leads to a worse objective value.
Fortunately, generating a feasible $L$-strong partition reduces to generating a feasible $1$-strong partition for a different ownership instance, as shown in the following proposition. This reduction helps us to convert any $L$-strongness constraint to a certain $1$-strongness constraint during our design of the optimization algorithm.
Proposition 1 (A Reduction from $L$-Strong to $1$-Strong Partition).
@@
For any ownership instance $\mathcal{O}$, we can construct in $O\left(m^{L}n\right)$ time a different ownership instance $\mathcal{O}^{\prime}$ such that any $L$-strong partition ${\mathcal{S}}$ in $\mathcal{O}$ is a 1-strong partition in $\mathcal{O}^{\prime}$.
@@
Proposition 1 shows that to optimize the informativeness objective, it is without loss of generality to design algorithms to maximize the objective under $1$-strongness, for a transformed ownership set instance. A simple example to illustrate this reduction: any $2$-strong partition in the ownership instance $\mathcal{O}=\{\{1,2\},\{1,2,3\},\{2,3\}\}$ is a $1$-strong partition in the ownership instance $\mathcal{O}^{\prime}=\{\{1,2\},\{2\},\{2,3\}\}$. Essentially, the instance $\mathcal{O}^{\prime}$ is constructed to have an owner with items shared by every $L$ owner in the original instance $\mathcal{O}$; we leave the details to the proof in Appendix. This is notably a polynomial reduction for any constant $L$ — e.g., the number of co-authors on each paper is typically small. However, this simplification does not fundamentally simplify the problem due to the intrinsic hardness of maximizing the informativeness objective, as shown in the following proposition.
@@
It is NP-hard to find the optimal item partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ that maximizes the size-based objective in Equation (2) subject to $1$-strongness.
@@ 4.2 Fast Greedy Partition with Robust Approximation Guarantees
Despite the computational hardness for the optimal partition, we show that a natural greedy algorithm (as also used in [23]) can efficiently find partitions with a provably good approximation guarantee. In words, the greedy algorithm just iteratively selects the largest residual paper set owned by some owner. This algorithm can be implemented in time almost linear in the input size $\sum_{j\in[m]}{|\mathcal{I}^{j}|}$, up to a $\log(m)$ factor, and thus is essentially the fastest algorithm one can design.
The highlight of this greedy algorithm is that, though the algorithm itself is fully agnostic to any partition objectives, its output is a robustly high-quality partition for every natural weight function, i.e., monotone and convex weight functions, as formalized in the following Theorem 2. In particular, consider the following hypothesis class $\mathcal{W}$ of weight functions
@@
Algorithm 3 enjoys the following strong approximation guarantees for every $w$ within $\mathcal{W}$.
@@
For any ownership instance $\{\mathcal{I}^{j}\}_{j\in[m]}$ with $N=\sum_{j\in[m]}{|\mathcal{I}^{j}|}$ total number of author-paper pairs, Algorithm 3 runs in $O(N\log m)$ time and outputs a 1-strong partition that is simultaneously a $c(w)$-approximation of the optimal objective for every $w\in\mathcal{W}$, where
@@
Input : Ownership sets $\{\mathcal{I}^{j}\}_{j\in[m]}$.
Initialize the partition as ${\mathcal{S}}=\{\}$ and set of selected items $\overline{\mathcal{I}}=\varnothing$.
while  ¯[n]$\overline{\mathcal{I}}\subset[n]$ do
       Determine the largest residual paper set ${\mathcal{S}}^{*}=\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$ where $j^{*}=\mathop{argmax}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$.
       Update ¯¯$\overline{\mathcal{I}}\leftarrow\overline{\mathcal{I}}\cup{\mathcal{S}}^{*}$ and ${\mathcal{S}}\leftarrow{\mathcal{S}}\cup\{{\mathcal{S}}^{*}\}$.
return ${\mathcal{S}}$.
Algorithm 3 A Greedy Algorithm for $1$-Strong Partition [23]
Remark 2.1.
@@
Theorem 2 readily implies that the partition found by the greedy algorithm simultaneously approximates at least $1/2$ of both the optimal comparison-based objective and size-based objective. More generally, for $\alpha$-th degree polynomial functions $w(x)=x^{\alpha}$, the greedy algorithm simultaneously guarantees $(1/\alpha)$-fraction of the optimal objective since $w(x)/[w^{\prime}(a)x]=1/\alpha$ for every $x$. Such objective-agnostic property of Algorithm 3 is especially desirable in our application scenarios where the concrete form of the partition objective is difficult (if not impossible) to know.
Moreover, the approximation ratio in Theorem 2 for the greedy algorithm is provably tight under every monomial objective $f({\mathcal{S}})=\sum_{k=1}^{K}|{\mathcal{S}}_{k}|^{\alpha}$, $\alpha>1$. Lower bound instances can be found in Section 6.2.1.
Remark 2.2.
While Algorithm 3 only finds a partition satisfying $1$-strongness, the algorithm as well as its constant approximation guarantees can be readily extended to find the optimal $L$-strong partition in $O\left(m^{L}n\right)$ time, using the reduction in Proposition 1.
5 Experiments
One major challenge in evaluating the proposed scoring mechanism is that their performance measures rely on the underlying true item qualities. Such information is unattainable in most applications — indeed, if we already know the true qualities, peer review would not be needed anymore. Fortunately, we are able to collect both the ownership relations and the review scores from ICLR 2021-2023 [1, 2, 3] that are made publicly available from openreview.net. Based on the recorded review scores, we simulate the true item qualities $\bm{r}=\textbf{y}-\bm{z}$, where $\bm{z}\sim\mathcal{N}(0,\sigma^{2})$ is a zero-mean Gaussian random variable with standard deviation $\sigma\in\{1,2,3,4\}$. We use the proposed Isotonic Mechanism to determine the calibrated score $\widehat{\bm{R}}$ from review scores y and the elicited ranking information. For comparison, we also consider the alternative calibration approaches: baseline, which directly uses the review scores; random, which runs the Isotonic Mechanism under a randomly generated partition. We measure the performance of the proposed mechanism by the mean squared error, $\frac{1}{n}\|\bm{R}-\widehat{\bm{R}}\|^{2}$, and the precision@kmetric, $\frac{|\widehat{\mathcal{P}}_{@k}\cap{\mathcal{P}}_{@k}|}{|{\mathcal{P}}_{@k}|}$, where ${\mathcal{P}}_{@k}$ (resp., $\widehat{\mathcal{P}}_{@k}$) is the set of items within $k$-th percentile with regard to the true quality $\bm{R}$ (resp., calibrated score $\widehat{\bm{R}}$). We choose the percentile $k\in\{1.5,5,30\}$ to roughly match with the criteria of “oral”, “spotlight” and “poster” presentations at ICLR, respectively.

### v3

@@ 4.2 Partition Optimization and Its Hardness
As illustrated above, any partition will make Mechanism 2 truthful and that partition-based Isotonic Mechanisms are the only truthful mechanisms within a much broader class of Isotonic Mechanisms. These together point to the last key piece of our design — identifying the item partition for the optimal statistical efficiency. In Remark 3, we make two key observations to formalize the relationship between partition structure and statistical efficiency.
Remark 3 (On Partition Optimization for Proximal Statistical Efficiency).
It is intuitive to see that both the number of owners and items of a partition block would influence the statistical performance of the review score calibration. We formalize these observations into two monotonicity laws about partition optimization. First, the coarser the partition is, the better the estimation accuracy. The proof is based on the observation that projecting onto a strictly smaller feasible region that contains the ground-truth leads to a smaller mean squared error in the estimated scores [37, Proposition 3.4]. Second, the more owners each partition block has, the better the estimation accuracy. This is straightforward when owners have independent perception noise about their item scores. That is, let $\widehat{\bm{R}}^{j}\sim\bm{R}+\bm{z}$ be the score estimate based on owner $j$’s reported ranking, where $\bm{R}$ is the true score vector and $\bm{z}$ consists of i.i.d. noise with bias $b$ and variance $\sigma^{2}$. Then, for any item $i$, its score estimate aggregated across $n$ owners, $\widehat{R}_{i}=\frac{1}{n}\sum_{i=1}^{n}\widehat{R}_{i}$, has mean squared error $b^{2}+\frac{\sigma^{2}}{n}$, which clearly decreases as the number of owners $n$ increases.
However, these monotonicity laws alone are not sufficient to solve the partition optimization problem. The set of feasible partitions is only partially ordered, and the rate at which noise is reduced remains difficult to quantify. As a result, precisely characterizing the relationship between partition structure and statistical efficiency remains an open question, if not an intractable one. Proving the optimality of statistical estimators is notoriously challenging, as exemplified by Stein’s paradox, where even the seemingly natural sample mean turns out to be suboptimal under certain conditions.
Following from Remark 3, we resort to the design of proxy objectives for partition optimization under two general principles: (1) Informativeness — larger blocks with more ranking information; [^12] For instance, the ranking of a size-$10$ item block has $\frac{10\times 9}{2}=45$ pairs of comparison, whereas the rankings of two size-$5$ item blocks have $202\times\frac{5\times 4}{2}=$ 20 pairs of comparison. (2) Robustness — each block has more owners. However, there is an inherent tension between the two principles, as larger blocks usually have fewer joint owners since it is less usual for multiple owners to jointly own many items. Thus the pursuit of larger blocks will prevent us from eliciting rankings from multiple owners. To address this tradeoff, we formalize two distinct metrics that quantify the informativeness and robustness of a partition and study the computational complexity of partition optimization for proximal statistical efficiency.
Informativeness (as optimization objective). In regard to the performance, we aim to identify partitions that are most informative for the review score calibration. On one hand, larger-sized blocks generally contain more ranking information, whereas in the extreme case if a block contains only one item, the mechanism cannot elicit any ranking information within the block. On the other hand, the selection of a large block may sometimes render other blocks small, leading to complex tradeoff about partition choices. To systemically evaluate the coverage quality of a partition $\mathfrak{S}$, we introduce an informativeness objective function,
$\operatorname{obj}(\mathfrak{S})=\sum_{k=1}^{K}w(|{\mathcal{S}}_{k}|),$
determined by some wellness function $w$ on the size of each partition block. For examples, consider the following two choices of the wellness function $w(\cdot)$. One is $w(x)=x^{2}$ as a simple quadratic function of block size, which relates to the number of pairwise comparisons within a block:
@@
Another choice is $w(x)=\mathop{max}\{x-1,0\}$, which is called size-focused objective, defined below:
@@
where the second equation is because any block in the partition has at least $1$ item (i.e., $|{\mathcal{S}}_{k}|\geq 1$). Since $n$ is a constant, maximizing the size-focused objective is equivalent to minimizing the size of the partition $K$ (thus the name “size-focused”).
Generally, it is reasonable to expect that function $w(\cdot)$ should be monotonically increasing and convex — to see convexity, if a block of length $l$ is broken into two blocks with length $l_{1},l_{2}$, then the original longer block must be preferred, i.e., $w(l)\geq w(l_{1})+w(l_{2})$ for any $l=l_{1}+l_{2}$, which is precisely the definition of convexity. Another natural requirement is $w(0)=0$ in order to prevent empty blocks to contribute any value to the partition wellness. Both functions above satisfy these properties, and so does any monomial $w(x)=x^{p}$ for $p\geq 1$. Beyond these properties, however, it appears impossible to exactly know the format of $w(\cdot)$. To overcome this challenge, we resort to robust algorithms in subsequent subsection 4.3 and look to design algorithms that can simultaneously perform well for every wellness function satisfying these properties.
@@
Robustness (as optimization constraint). If the informativeness metric above is regarded as the objective we want the partition to maximize, then the robustness metric can be viewed as a constraint we wish the partition to satisfy. In regard to robustness, aggregating information from multiple owners can help mitigate each individual owner’s noisy perception of their items in situations where Assumption 2.1 does not hold. Thus, the minimal number of owners for each partition block matters. Formally, we say a partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ is $L$-strong if there are at least $L$ owners who share all items in the same partition block, i.e., $\left|{\mathcal{T}}_{k}\right|\geq L,\forall{\mathcal{S}}_{k}\in\mathfrak{S}$ with $|{\mathcal{S}}_{k}|>1$. The larger $L$ is, the less prone Isotonic Mechanism is to the noise in each owner’s perceived ranking. Notably, $L\geq 1$ is the bare minimum requirement for Mechanism 2 to work. Next we show that generating a feasible $L$-strong partition reduces to generating a feasible $1$-strong partition for a different ownership instance. This reduction helps us convert any $L$-strongness constraint to a certain $1$-strongness constraint during our design of the partition optimization algorithm.
Proposition 1 (Reduction to $1$-Strong Partition).
@@
For any ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$, we can construct in $O\left(m^{L}n\right)$ time a different ownership instance $\mathcal{O}^{\prime}=\{\overline{\mathcal{I}}^{j}\}_{j\in[m]}$ with the same owner and item set such that any partition ${\mathcal{S}}$ is $L$-strong in $\mathcal{O}$ if and only if ${\mathcal{S}}$ is $1$-strong in $\mathcal{O}^{\prime}$.
@@
Proposition 1 shows that to optimize the informativeness objective, it is without loss of generality to design algorithms to maximize the objective under $1$-strongness, for a transformed ownership set instance. We defer its formal proof to Appendix B but provide a simple example below to illustrate the main idea of the reduction: any $2$-strong partition in the ownership instance $\mathcal{O}=\{\{1,2\},\{1,2,3\},\{2,3\}\}$ is a $1$-strong partition in the ownership instance $\mathcal{O}^{\prime}=\{\{1,2\},\{2\},\{2,3\}\}$, and vice versa. Essentially, the instance $\mathcal{O}^{\prime}$ is constructed by using one owner to represent each subset of $L$ owners in the original instance $\mathcal{O}$, whose common items are owned by this owner. This is a polynomial-time reduction for small constant $L$, which is often the case in academic conference review due to the small number of co-authors on a paper. Notably, this reduction does not fundamentally simplify the problem due to the intrinsic hardness of maximizing the informativeness objective, as shown in the following proposition; see Appendix C for its formal proof.
@@
It is NP-hard to find the optimal partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ that maximizes the size-focused objective in Eq. (3) subject to $1$-strongness.
@@ 4.3 Fast Greedy Partition with Robust Approximation Guarantees
Input: Ownership sets $\{\mathcal{I}^{j}\}_{j=1}^{m}$.
Initialize the partition as $\mathfrak{S}=\{\}$ and set of selected items $\overline{\mathcal{I}}=\varnothing$.
while  ¯$\overline{\mathcal{I}}\subset[n]$ do
       Determine the largest residual item set ${\mathcal{S}}^{*}=\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$ where $j^{*}=\mathop{argmax}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$.
       Update ¯¯$\overline{\mathcal{I}}\leftarrow\overline{\mathcal{I}}\cup{\mathcal{S}}^{*}$ and $\mathfrak{S}\leftarrow\mathfrak{S}\cup\{{\mathcal{S}}^{*}\}$.
return $\mathfrak{S}$.
Algorithm 4 A Greedy Algorithm for $1$-Strong Partition
Despite the computational hardness for the optimal partition, we show that a natural greedy algorithm presented in Algorithm 4 can efficiently find $1$-strong partitions with a provably good approximation guarantee. In words, this greedy algorithm just iteratively selects the largest residual item set owned by some owner. This algorithm can be implemented in time almost linear in the input size $\sum_{j\in[m]}{|\mathcal{I}^{j}|}$, up to a $\log(m)$ factor, thus is essentially the fastest algorithm one could hope to design.
The highlight of this greedy algorithm is that, though the algorithm itself is fully agnostic to any partition objectives, its output is a robustly high-quality partition for every natural wellness function, i.e., monotone and convex wellness functions, as formalized in the following Theorem 3. In particular, consider the following hypothesis class $\mathcal{W}$ of wellness functions
@@
Algorithm 4 enjoys the following strong approximation guarantees for every $w$ within $\mathcal{W}$, whose proof is deferred to Appendix E. In Remark 4, we give the exact approximation ratio for some optimization objectives of special interests and show its tightness. It is also worth mentioning that while Algorithm 4 only finds a partition satisfying $1$-strongness, the algorithm as well as its constant approximation guarantees can be readily extended to find the optimal $L$-strong partition in $O\left(m^{L}n\right)$ time, using the reduction in Proposition 1.
@@
For any ownership instance $\{\mathcal{I}^{j}\}_{j\in[m]}$ with $N=\sum_{j\in[m]}{|\mathcal{I}^{j}|}$ total number of owner-item pairs, Algorithm 4 runs in $O(N\log m)$ time and outputs a 1-strong partition that is simultaneously a $c(w)$-approximation of the optimal objective for every $w\in\mathcal{W}$, where
@@
Remark 4.
@@
Theorem 3 readily implies that the partition found by the greedy algorithm simultaneously approximates at least $1/2$ of both the optimal comparison-focused objective and size-focused objective. More generally, for $\alpha$-th degree polynomial functions $w(x)=x^{\alpha}$, the greedy algorithm simultaneously guarantees $(1/\alpha)$-fraction of the optimal objective since $w(x)/[w^{\prime}(a)x]=1/\alpha$ for every $x$. Such objective-agnostic property of Algorithm 4 is especially desirable in our application scenarios where the concrete form of the partition objective is difficult (if not impossible) to know. Moreover, the approximation ratio in Theorem 3 for the greedy algorithm is provably tight under every monomial objective $f(\mathfrak{S})=\sum_{k=1}^{K}|{\mathcal{S}}_{k}|^{\alpha}$, $\alpha>1$, and the construction of lower bound instances can be found in Appendix E.1.
5 Empirical Evaluation in the Context of ML Conference Review
Figure 4: An illustration of applying our proposed Isotonic Mechanism to the conference reviewing procedure.
In this section, we study the empirical performance of our proposed mechanism by applying it to our primary motivating domain of academic conference peer review. An illustrative procedure of applying our mechanism to conference peer review can be found in Figure 4. One major challenge of evaluation is that its performance measures rely on the underlying ground-truth scores, and such information is unattainable in most applications — indeed, if we already know the ground-truth scores, peer review would not be needed anymore. Fortunately, from both ICLR and ICML, we are able to collect some parts of their data and synthesize some of the unobservable parts. Below, we describe our experiment setups and results.
5.1 Experiment Setups: Datasets, Baselines and Metrics
We start by elaborating on the preprocessing procedures and characteristics of the two real-world datasets we experimented on.
  1. 1.
    ICLR 2021-2023: We collect ownership relations and review scores from ICLR 2021-2023 [17, 18, 19] that are made publicly available from OpenReview.net. Based on the recorded review scores, we simulate the ground-truth scores $\bm{R}=\textbf{y}-\bm{z}$, where $\bm{z}\sim\mathcal{N}(0,\sigma^{2})$ is a zero-mean Gaussian random variable with standard deviation $\sigma\in\{1,2,3,4\}$.
  2. 2.
    ICML 2023: We collect ownership relations, review scores as well as the authors’ reported ranking on their papers from ICML 2023, based on survey from OpenRank.cc. We only keep the authors (and their papers) who have participated in the survey and provided a ranking of their submissions. For each paper, we split its review scores by their reviewers’ confidence level. We set the review score with the least confidence level as the paper’s raw review score $\bm{y}$, and the average score of remaining reviews as the paper’s ground-truth score $\bm{R}$.
Notably, we adopted two different approaches to generate the ground-truth scores in the two datasets above. While neither approach can perfectly mimic the ground-truth score, they help us understand the effectiveness of the Isotonic Mechanism from different perspectives. For the ICLR dataset, we generate the ground-truth scores based on the i.i.d. review score noises assumption. In this way, we are able to empirically test the Isotonic Mechanism under the score generation model of this paper. For the ICML dataset, we use the high confidence review scores (most likely from domain experts) as the surrogate to approximate the ground-truth scores. While this score generation process does not necessarily match our model, it allows us to directly observe whether the Isotonic Mechanism can adjust review scores with high noise towards the consensus of experts, which tends to better align with the ground-truth score.
We use the Mechanism 2 with uniform owner credential ($\alpha^{j}=1,\forall j$) to determine the adjusted score $\widehat{\bm{R}}$ from review scores y and the elicited ranking information. For comparison, we consider the following alternative calibration approaches, evaluated under the mean squared error (MSE), $\frac{1}{n}\|\bm{R}-\widehat{\bm{R}}\|^{2}$: baseline, which directly uses the raw review scores; random, which uses a randomly generated partition for the Isotonic Mechanism.
Table 4 summarizes the basic statistics of ICLR 2021-2023 that we collected as well as the partitions determined from their ownership sets. There is a clear growing trend in conference size in terms of submission and author number. Moreover, from both partition objective functions, the greedy partition is a clearly better choice than the randomly generated partition, though we are unable to empirically examine the approximation ratio of Greedy (Algorithm 4), since it is intractable to find the exactly optimal partition for such large instances given its NP-hardness.